OpenStudy (anonymous):

how can i solve this problem? please help me.. i've been stuck on it for days now !

OpenStudy (anonymous):

do you know what g(x) is?

OpenStudy (mathmale):

Euler's question is a good one. You can (or should be able to) obtain the function g(x) just from looking at the blue graph. What is the slope? What is the y-intercept?

zepdrix (zepdrix):

Ooo fun problem :)

zepdrix (zepdrix):

I guess finding your g(x) might be a good idea. I would just use the chain rule and the slope of g(x) though. \[\Large\bf\sf w(x)=g(g(x))\]\[\Large\bf\sf w'(x)=?\]Do you understand how to take this derivative?

OpenStudy (mathmale):

Again, that blue line represents the function g(x), and again I suggest that you obtain that function before finding any derivatives.

OpenStudy (anonymous):

so for a would you plug in 4 for x? ... so g(g(4)) ... but the derivative?

zepdrix (zepdrix):

You can't plug 4 in until you've differentiated first. Do you understand how to differentiate g(g(x))?

OpenStudy (anonymous):

no

OpenStudy (mathmale):

@skullofreak: I've twice suggested that you obtain g(x) from looking at the graph. Then, you need to differentiate g(g(x)).

zepdrix (zepdrix):

Here's the definition of the chain rule to refresh your memory:\[\Large\bf\sf y=f(g(x)), \qquad y'=f'(g(x)\cdot g'(x)\]

OpenStudy (mathmale):

You decide. If you want my advice, please determine g(x) from the graph.

zepdrix (zepdrix):

I don't particularly like that method. It's easier yes, but you really need to get comfortable with the chain rule.

zepdrix (zepdrix):

It's one of the most confusing differentiation rules when you're early in calc, you really need to spend a lot of time with it :[

zepdrix (zepdrix):

You take the derivative of the outer function, then multiply by the derivative of the inner function. \[\Large\bf\sf y'=f'(g(x))\cdot g'(x)\]

zepdrix (zepdrix):

What's confusing? the fact that we're not given an actual function? It is it not specific enough so it's giving you trouble?

OpenStudy (anonymous):

so i look at 4 then find the values for f(x) and g(x)? or how would i get the values/..

zepdrix (zepdrix):

That was simply a definition. In our case we have g(g(x)), not f(g(x)), so it will look a little different. \[\Large\bf\sf w(x)=g(g(x))\]Using the definition for the chain rule,\[\Large\bf\sf w'(x)=g'(g(x))\cdot g'(x)\]You can plug in your 4 from here,\[\Large\bf\sf w'(4)=g'(g(4))\cdot g'(4)\]And then use the graph to fill in the pieces. If this is way too confusing, then use male's method. It will give you an actual function that you can differentiate instead of reading a graph.

zepdrix (zepdrix):

Reading the graph,\[\Large\bf\sf g(4)=12\]So we'll plug that in,\[\Large\bf\sf w'(4)=g'(\color{orangered}{g(4)})\cdot g'(4)\]\[\Large\bf\sf w'(4)=g'(\color{orangered}{12})\cdot g'(4)\]

OpenStudy (anonymous):

so 8(12(4))*8(4)?.

OpenStudy (anonymous):

Ohh nvm

OpenStudy (anonymous):

wait would derivative of g be 8 then? cause it is f(x)? or that has nothing to do with it

zepdrix (zepdrix):

Hmm maybe I'm misunderstanding the question... This doesn't appear to have ANYTHING to do with the f(x) that was graphed. I can't figure out why it's there. :(

zepdrix (zepdrix):

Is there a typo in the problem? I think it's supposed to be w(x) = f(g(x)) or something like that... it would make a lot more sense.

OpenStudy (anonymous):

hmm well theres 5 more problems like this and one has f(f(x)) too ..

zepdrix (zepdrix):

Oh we use this same picture for a bunch of different problems? Ok then it seems fine :)

zepdrix (zepdrix):

We're just not using the f for any part of this.

OpenStudy (anonymous):

no the graph is just for A,B and C the other questions have different graphs

zepdrix (zepdrix):

o

OpenStudy (anonymous):

like this one is f(f(x))

zepdrix (zepdrix):

Ya I guess they do this sometimes in math problems. They give you more information than you actually need.

zepdrix (zepdrix):

The problem simplified to, \[\Large\bf\sf w'(4)=g'(12)\cdot g'(4)\] Now you need the slope of g at x=12, and the slope at x=4. Since it's a linear function, both values will be the same (constant slope).

OpenStudy (anonymous):

the derivative of g would be 8 then?

zepdrix (zepdrix):

No. Hmm.. I'm not sure where you're getting 8 from. See how the line slants `down` to the right? It has a negative slope.

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Do you remember how to find the slope of a line.?

OpenStudy (anonymous):

rise/run

zepdrix (zepdrix):

Good. So use a couple of points to find the slope. I would recommend using the (16,0) and (0,16).

OpenStudy (anonymous):

so... y2-y1/x2-x1 = 16-0/0-16

OpenStudy (anonymous):

= -1

zepdrix (zepdrix):

Yes, good!

zepdrix (zepdrix):

g'(12) is the `slope` of g(x) at x=12. And we're saying... that since it's a straight line, it has a slope of -1 everywhere.

zepdrix (zepdrix):

Which means g'(12) = -1 and also g'(4) = -1

OpenStudy (anonymous):

so the slope we get would be the derivative?

OpenStudy (anonymous):

so there isnt a slope for the derivative of 8, but when i put DNE it says its wrong

OpenStudy (anonymous):

because of the 'sharp corner'?

zepdrix (zepdrix):

There is no sharp corner for the function g(x)

OpenStudy (anonymous):

so a,b , and c would just be -1?

zepdrix (zepdrix):

No none of them are -1.

OpenStudy (anonymous):

oh yeah it was g(g(x)) we arent using f (x)

zepdrix (zepdrix):

\[\Large\bf\sf w'(4)=g'(12)\cdot g'(4)\]\[\Large\bf\sf w'(4)=(-1)\cdot (-1)\]

OpenStudy (anonymous):

oh ! okay okay i get it

zepdrix (zepdrix):

But ya it looks like they all work out to the same value for a b and c.

OpenStudy (anonymous):

yes b would be g'(8)*g'(8) = 1 again , ty so much! :) i can do the rest now !